What is the formula to find out if the equations are parallel, perpendicular, or neither? y = 3x – 1, y = –x + 2
Answer by Alan3354(69443)

What is the formula to find out if the equations are parallel, perpendicular, or neither? y = 3x – 1, y = –x + 2
------------
There's no formula, there's a method.
Find the slope of each line, m.
if the slopes are:
equal they're parallel
negative reciprocals they're perpendicular
anything else, neither.
----------------------
y = 3x – 1 slope = 3
y = –x + 2 slope = -1
These are not parallel and not perpendicular.